OpenStudy (studytheworld):

Please help! Will medal and fan! Helppp! In the triangle below, b = ____. If necessary, round your answer to two decimal places. (See link)

OpenStudy (studytheworld):

http://prntscr.com/8zhasx

OpenStudy (studytheworld):

@TrojanPoem You know this one?

OpenStudy (trojanpoem):

A + B + C = 180 B = 180 - 48 - 46.7 B = 85.3 b/sinb = a/sina b = a sinb/sina 44 * sin(85.3)/sin46.7 = 60.26

OpenStudy (studytheworld):

B = 60.26?

OpenStudy (studytheworld):

or 85.3?

OpenStudy (trojanpoem):

In triangles we refer to the angle with capital letters , small letters for the side lengths.

OpenStudy (trojanpoem):

The angles of the triangle sum up to 180 , so A + B + C = 180

OpenStudy (studytheworld):

Okay

OpenStudy (studytheworld):

And?

OpenStudy (trojanpoem):

And we use the normal sine law.

OpenStudy (studytheworld):

Okay so is my answer 60.26?

OpenStudy (trojanpoem):

Yeah, if I didn't miscalculate a number xD

OpenStudy (studytheworld):

Thanks so much for helping me